# PhysBrain: Human Egocentric Data as a Bridge from Vision Language Models to Physical Intelligence

### v1

@@ 5.2 VLA Simulation Evaluation
To validate the efficacy of our model when deployed as the VLA for robotic control, we adopt PhysBrain as the VLM backbone and fine-tune it within the VLA paradigm using downstream robotics data. We then evaluate on the SimplerEnv (Li et al., 2024c) simulation benchmark with the WidowX robot.
@@ 5.2.1 Experiment Settings
Architecture. We instantiate the VLA model using the PhysGR00T and PhysPI architecture as described in Sec. 4. The VLM component is initialized with weights from PhysBrain, whereas the Action Expert is initialized with random weights.
[…]
Evaluation. The benchmark consists of four manipulation tasks: "put spoon on towel", "put carrot on plate", "stack green block on yellow block", "put eggplant in the yellow basket". For each task, we evaluate our VLA policy using the official evaluation script provided by the SimplerEnv repository (Li et al., 2024c). To mitigate randomness, we run five independent trials and report the mean performance.
Baselines. We primarily compare our method against two categories of baselines: (i) VLA baselines, which include several widely used VLA models (RT-1-X, Octo, OpenVLA, RoboVLM, TraceVLA, SpatialVLA, CogACT, VideoVLA and $\pi_{0}$); and (ii) VLM baselines, where we fine-tune several commonly used VLMs (RoboBrain2.0, VST-RL and Spatial-SSRL) under the VLA paradigm and evaluate them using the same training configuration as our method.
Table 2: Results of evaluating the VLA models with the WidowX robot in the SimplerEnv simulation environment, where the VLM backbone is fine-tuned under the VLA paradigm following the PhysGR00T architecture. We highlight the best results in bold and the second-best results with underline.
Method Put Spoon on Towel Put Carrot on Plate Stack Green Block on Yellow Block Put Eggplant in Yellow Basket Average VLA Baselines RT-1-X (O’Neill et al., 2024) 0.0 4.2 0.0 0.0 1.1 Octo-Base (Team et al., 2024) 15.8 12.5 0.0 41.7 17.5 Octo-Small (Team et al., 2024) 41.7 8.2 0.0 56.7 26.7 OpenVLA (Kim et al., 2024) 4.2 0.0 0.0 12.5 4.2 OpenVLA-OFT (Kim et al., 2025) 12.5 4.2 4.2 72.5 23.4 RoboVLM (Li et al., 2024b) 50.0 37.5 0.0 83.3 42.7 TraceVLA (Zheng et al., 2025) 12.5 16.6 16.6 65.0 27.7 SpatialVLA (Qu et al., 2025) 20.8 20.8 25.0 70.8 34.4 CogACT (Li et al., 2024a) 71.7 50.8 15.0 67.5 51.3 VideoVLA (Shen et al., 2025) 75.0 20.8 45.8 70.8 53.1 $\pi_{0}$ (Black et al., 2024) 29.1 0.0 16.6 62.5 27.1 $\pi_{0}$-FAST (Pertsch et al., 2025) 29.1 21.9 10.8 66.6 48.3 VLM Baselines Qwen2.5-VL-7B (Bai et al., 2025b) 59.2 30.8 3.3 44.2 34.4 RoboBrain2.0-7B (Team et al., 2025) 30.8 24.7 2.5 93.3 37.8 VST-RL-7B (Yang et al., 2025a) 57.7 41.7 16.7 50.0 41.3 Spatial-SSRL-7B (Liu et al., 2025) 56.3 44.8 6.2 72.9 45.1 PhysBrain (ours) 65.6 37.5 33.3 79.2 53.9
@@ 5.2.2 Experiment Results
Table 2 summarizes the SimplerEnv evaluation results, comparing our PhysBrain model, fine-tuned under the VLA paradigm following the PhysGR00T architecture, against all baseline methods. More evaluation results under the PhysPI architecture are presented in Appendix 8.
(i) Comparison with VLA Baselines. Despite being fine-tuned on only two subsets of the OXE dataset (Bridge and Fractal), our VLA model achieves an average success rate of 53.9%, outperforming VLA baselines trained on substantially larger robot datasets (e.g., the full OXE dataset comprising 55 subsets). This improvement demonstrates that egocentric human data, when properly annotated and leveraged during pretraining, can effectively compensate for the robot-specific data.
(ii) Comparison with VLM Baselines. Under the same training paradigm, we fine-tune several commonly used open-source VLMs into VLA models for comparison. As demonstrated in Table 2, our model consistently outperforms all VLM baselines across all tasks, achieving an average improvement of 8.8% over the second-best performing model and a substantial 16.1% gain over RoboBrain (Team et al., 2025), which is specifically designed for embodied intelligence tasks. These results provide evidence that VLMs pretrained on large-scale human egocentric data yield more effective initialization for downstream VLA fine-tuning. Notably, the domain-agnostic generalization capabilities induced by egocentric human data enable successful VLA training with only a limited amount of robot-specific data, highlighting the transferability of human behavioral priors to robotic manipulation.
@@ 6 Conclusion
In this work, we address the fundamental challenge of exploiting human egocentric videos to bridge vision-language models with physical intelligence for robotic generalization. We introduce an Egocentric2Embodiment translation pipeline that systematically converts raw human egocentric videos into multi-level, schema-driven VQA supervision with deterministic rule validation, producing the E2E-3M dataset with approximately three million verified instances across household, factory, and laboratory domains. By supervised fine-tuning on this dataset without requiring any robot-collected data for VLM pretraining, we develop PhysBrain, which substantially improves egocentric capability (particularly in planning) as demonstrated by the EgoThink Benchmark, and achieves high success rate on SimplerEnv when used as the VLM backbone in standard VLA fine-tuning. Our results validate that scalable human egocentric supervision can serve as a practical and effective bridge from vision-language understanding to physical intelligence, opening promising directions for expanding egocentric data diversity, developing more sophisticated translation mechanisms, and exploring efficient policy learning from human demonstrations.
Limitations and Future Work
While our work demonstrates that VLMs pretrained on human egocentric data yield effective pretrained checkpoints for VLA training, several limitations warrant further investigation. First, our experimental evaluation primarily focuses on the PhysGR00T architecture with limited exploration of the PhysPI variant. A more comprehensive analysis encompassing diverse architectural configurations, refined experimental protocols, and systematic ablation studies remains to be conducted. Second, further investigations into the complementarity between human egocentric data and robot demonstration data are ongoing. We plan to progressively release these additional experimental results and extended analyses in subsequent versions of this work.
Acknowledgements
This work is supported by the Zhongguancun Institute of Artificial Intelligence (Grant No.A20250501). Additionally, we would like to extend our special gratitude to DeepCybo for their assistance with data annotation and other processes.

### v2

@@ 5.2 VLA Simulation Evaluation
Table 3: Results of evaluating the VLA models with the GR1 robot in the RoboCasa Tabletop simulation environment. The results for QwenGR00T, QwenOFT, and QwenFAST are derived from the official StarVLA experiments (starVLA Community, 2025). We highlight the best results in bold and the second-best results with underline.
Task Isaac-GR00T N1.6 QwenGR00T + Qwen3VL QwenOFT + Qwen3VL QwenFAST + Qwen3VL PhysBrain-4B PhysBrain-8B
PnP Bottle To Cabinet Close 51.5 46.0 30.0 38.0 74.0 70.0
PnP Can To Drawer Close 13.0 80.0 76.0 44.0 68.0 74.0
PnP Cup To Drawer Close 8.5 54.0 44.0 56.0 42.0 46.0
PnP Milk To Microwave Close 14.0 48.0 44.0 44.0 54.0 60.0
PnP Potato To Microwave Close 41.5 28.0 32.0 14.0 24.0 34.0
PnP Wine To Cabinet Close 16.5 46.0 36.0 14.0 54.0 40.0
PnP Novel From Cuttingboard To Basket 58.0 48.0 50.0 54.0 62.0 54.0
PnP Novel From Cuttingboard To Cardboardbox 46.5 40.0 40.0 42.0 44.0 56.0
PnP Novel From Cuttingboard To Pan 68.5 68.0 70.0 58.0 56.0 72.0
PnP Novel From Cuttingboard To Pot 65.0 52.0 54.0 58.0 58.0 74.0
PnP Novel From Cuttingboard To Tieredbasket 46.5 56.0 38.0 40.0 40.0 44.0
PnP Novel From Placemat To Basket 58.5 42.0 32.0 36.0 42.0 58.0
PnP Novel From Placemat To Bowl 57.5 44.0 58.0 38.0 56.0 56.0
PnP Novel From Placemat To Plate 63.0 48.0 52.0 42.0 80.0 62.0
PnP Novel From Placemat To Tieredshelf 28.5 18.0 24.0 18.0 14.0 28.0
PnP Novel From Plate To Bowl 57.0 60.0 60.0 52.0 54.0 70.0
PnP Novel From Plate To Cardboardbox 43.5 50.0 50.0 30.0 50.0 54.0
PnP Novel From Plate To Pan 51.0 54.0 66.0 48.0 68.0 56.0
PnP Novel From Plate To Plate 78.7 70.0 68.0 50.0 78.0 60.0
PnP Novel From Tray To Cardboardbox 51.5 38.0 44.0 28.0 40.0 52.0
PnP Novel From Tray To Plate 71.0 56.0 56.0 34.0 66.0 60.0
PnP Novel From Tray To Pot 64.5 50.0 62.0 46.0 52.0 70.0
PnP Novel From Tray To Tieredbasket 57.0 36.0 54.0 36.0 50.0 48.0
PnP Novel From Tray To Tieredshelf 31.5 16.0 30.0 16.0 22.0 28.0
Average 47.6 47.8 48.8 39.0 49.75 55.25
To validate the efficacy of our model when deployed as the VLA for robotic control, we adopt PhysBrain as the VLM backbone and fine-tune it within the VLA paradigm using downstream robotics data. We then evaluate on the SimplerEnv (Li et al., 2024c) and RoboCasa (Nasiriany et al., 2024) simulation benchmarks.
SimplerEnv: The WidowX Robot is utilized in the SIMPLER environment to evaluate the VLA policy on four manipulation tasks: "Put Spoon on Towel," "Put Carrot on Plate," "Stack Green Block on Yellow Block," and "Put Eggplant in Yellow Basket." For each task, we conduct 24 trials and using the official evaluation script from the SimplerEnv repository. This ensures a consistent and rigorous assessment of the robot’s manipulation performance across these tasks.
RoboCasa: RoboCasa GR1 Tabletop Manipulation Benchmark includes 24 diverse tasks involving complex interactions with articulated objects and varied geometries. Examples of tasks include "PnPBottleToCabinetClose" and "PnPCanToDrawerClose", as well as scenarios with appliances like microwaves and toasters. For training, we use the Humanoid Robot Tabletop Manipulation subset from the PhysicalAI-Robotics-GR00T-X-Embodiment-Sim dataset (Bjorck et al., 2025b), ensuring a robust and diverse training environment for evaluating robotic manipulation performance.
@@ 5.2.1 Experiment Settings
Training. To adapt the VLM to the VLA architecture and the target robotic platform, we follow the training configuration of the starVLA (starVLA Community, 2025) framework. Detailed training hyperparameters and configurations are provided in Appendix 9.
Baselines. We primarily compare our method against two categories of baselines: (i) VLA baselines, which include several widely used VLA models (RT-1-X, Octo, OpenVLA, RoboVLM, TraceVLA, SpatialVLA, CogACT, VideoVLA, $\pi_{0}$, $\pi_{0.5}$, and Issac-GR00T-N1.6-Bridge); and (ii) VLM baselines, where we fine-tune several commonly used VLMs (Qwen2.5-VL, Qwen3-VL, RoboBrain2.0, VST-RL, and RoboBrain2.5) under the PhysVLA paradigm and evaluate by using the same training configuration as ours.
Evaluation. We evaluate our policy on SimplerEnv (Li et al., 2024c) (four WidowX tasks) and RoboCasa (24 complex manipulation tasks involving articulated objects). To ensure statistical robustness, we adhere to official protocols, reporting mean success rates (Avg@50) averaged over 5 and 50 independent trials for SimplerEnv and RoboCasa, respectively.
@@ 5.2.2 Experiment Results
Table 2 summarizes the SimplerEnv evaluation results, comparing our PhysBrain with all baseline methods. The evaluation results of RoboCasa are presented in Table 3.
As detailed in Table 2, PhysBrain-8B achieves an average success rate of 67.4%, significantly surpassing VLA baselines trained on substantially larger robot datasets (e.g., Isaac-GR00T at 57.1%) and performing comparably to the state-of-the-art RoboBrain2.5 (67.6%). Crucially, while RoboBrain2.5 relies on massive cross-embodiment robot data for representation alignment, PhysBrain attains this proficiency solely through pretraining on our proposed E2E-3M dataset. This highlights the potential of human data by providing robust physical priors that transfer seamlessly to robotic manipulation.
We evaluate PhysBrain on the RoboCasa benchmark against the Qwen3-VL-4B backbone integrated with three distinct action encodings (GR00T, OFT, and FAST), as shown in Table 3. Since PhysBrain adopts the Gr00T architecture, the comparison with the standard QwenGR00T baseline is particularly critical, as it directly isolates the performance gains attributable to our proposed method. As shown in Table 3, PhysBrain-4B achieves a 49.75% success rate, not only surpassing its direct counterpart (47.8%) but also outperforming other encoding variants. Furthermore, the PhysBrain-8B model validates significant scalability, attaining the best average success rate of 55.25%.
5.3 Ablation Study
Table 4: Ablation experiment on the impact of the proposed E2E-3M dataset across VLM Architectures and Model Scales. Best results are in bold.
Method VLM Bench VLA Bench
Qwen2.5-VL-7B 58.7 34.4
PhysBrain2.5-7B (ours) 63.1 53.9
Qwen3-VL-4B 66.7 55.2
PhysBrain-4B (ours) 69.4 65.9
Qwen3-VL-8B 65.9 56.3
PhysBrain-8B (ours) 69.7 67.4
Dataset Impact Across VLM Architectures and Model Scales. To verify the universality of the proposed E2E-3M dataset, we conducted extensive ablation studies across different VLM architectures (Qwen2.5-VL and Qwen3-VL) and varying parameter scales (4B, 7B, and 8B). As detailed in Table 4, fine-tuning on our dataset consistently yields significant performance improvements over the base models across all configurations. Notably, while the VLM benchmark scores show steady increments, the improvements on the VLA benchmark are particularly substantial (e.g., a 19.5% gain for the 7B model). This disparity indicates that while standard VLMs possess strong general vision-language capabilities, they lack the specific physical priors required for robotic manipulation. These results confirm that the physical intelligence distilled from E2E-3M is agnostic to the underlying architecture, consistently enhancing the embodied planning capabilities of diverse foundation models.
Table 5: Ablation experiment on the impact of the egocentric embodied supervision scale. Best results are in bold.
Method VLM Bench VLA Bench
PhysBrain-8B-wo-Ego4D 67.8 64.1
PhysBrain-8B (ours) 69.7 67.4
Impact of Egocentric Supervision Data Scale. We further investigate the scaling behavior of our approach by analyzing the impact of dataset size on model performance. Table 5 presents an ablation study where we remove the large-scale Ego4D subset from the training data. The results demonstrate a clear positive correlation between the volume of egocentric supervision and the resulting model proficiency. The full PhysBrain-8B model outperforms the version trained without Ego4D on both VLM and VLA benchmarks. This decline in performance upon data reduction highlights that the scale and diversity of the dataset are critical factors. It suggests that scaling up high-quality egocentric human data is a viable and effective pathway for continuously improving the physical understanding and manipulation skills of embodied agents.
@@ 6 Conclusion
In this work, we leverage human egocentric videos to connect vision-language models with physical intelligence for robotic generalization. We introduce the Egocentric2Embodiment pipeline, which converts raw egocentric videos into multi-level VQA supervision, creating the E2E-3M dataset with nearly three million verified instances across household, factory, and laboratory settings. By fine-tuning a vision-language model on the E2E-3M dataset, we develop PhysBrain that significantly enhances egocentric capabilities. It also achieves high success rates on VLA system when used as the VLM backbone in standard VLA fine-tuning. Our results demonstrate that scalable human egocentric supervision bridges vision-language understanding and physical intelligence, enabling greater data diversity and improved policy learning from human demonstrations.